An ideal refrigerator-a Carnot engine running backwards-freezes ice-cubes at a rate of 5gs-1 starting with water at its freezing point. Energy is given off to the room which is at 30°C. The fusion energy of ice is 320Jg-1. At what rate must electrical energy be supplied as work, and at what rate is energy given off to the room
1
Expert's answer
2019-04-05T10:21:54-0400

The rate energy is given off to the room equals to Pr = Pf/(1-η) = mλTh/Tl = 5*320*303/273 = 1.8 (kW). The electrical energy rate Pel = Pr – Pf = 1800 – 5*320 = 200 (W).